A drama class has a total 24 students the number of males is 12 more than the number of females how many males and how many females are in a class
Answer by math_tutor2020(3816)


x = number of women in the class
x+12 = number of men in the class

Add up the two amounts, set them equal to 24. Solve for x.
women + men = 24
x + (x+12) = 24
2x+12 = 24
2x = 24-12
2x = 12
x = 12/2
x = 6

Or you can solve like this
x + (x+12) = 24
2x+12 = 24
2(x+6) = 24
x+6 = 24/2
x+6 = 12
x = 12-6
x = 6

x+12 = 6+12 = 18

Answers:
6 women and 18 men